OpenStudy (anonymous):

What is the slope of the line 5x + 3y = 15?

OpenStudy (anonymous):

its 5x right?

OpenStudy (anonymous):

5x-3y=15 -3y=-5x+15 y=5/3x-5 y=mx+b m is slope b is y-intercept the slope is 5/3 because I subtract 5x by both sides and divide -3 both sides.

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yeah, thank you (:

OpenStudy (anonymous):

No problem! :D

OpenStudy (anonymous):

It's -5/3 -_-